# MARVIS: Motion & Geometry Aware Real and Virtual Image Segmentation

### v1

@@ I INTRODUCTION
Advanced computer vision systems enable robots to comprehend scenes by analyzing pictorial information. Although this may work well in many cases, various real world scenarios can pose challenges due to the diverse nature of light propagation and scene complexity. The images received by human and machine visual sensors are often a complex superposition of direct imaging, reflections, refractions, scattering, and the likes. This complexity is particularly pronounced at the interfaces between different media. Consequently, the input to vision sensors contains real image regions formed by direct imaging and virtual areas typically formed by distortion, reflection, or refraction. Inherent complexities in differentiating between real and virtual regions pose challenges for both low-level and high-level computer vision tasks (see Fig. 1), such as depth and pose estimations, visual odometry, object detection and recognition, semantic segmentation, etc. Therefore, real-virtual image segmentation to localize visually credible areas is crucial for marine robots, particularly in multimedia scenarios.
@@
Novel deep learning-based methods have significantly improved the performance of numerous computer vision tasks such as object recognition [1] and detection [2], 3D navigation [3], and action understanding [4]. However, our investigation indicates that real-virtual image segmentation in multimedia scenarios remains a relatively underexplored area. One primary challenge is the scarcity of publicly available datasets for real-virtual image segmentation in multimedia scenarios. Additionally, existing underwater simulators often overlook the intricate modeling of water-air interface imaging, making synthetic data extremely difficult to acquire. At the same time, data-driven methods are heavily relying on large-scale datasets. Secondly, the absence of an efficient annotation pipeline for the real-virtual image segmentation task, particularly for images with visual scene ambiguity, results in manual annotation being more laborious than common computer vision tasks. Moreover, achieving this task usually involves the use of Unmanned Aerial Vehicles (UAVs), Autonomous Surface Vehicles (ASVs), Unmanned Underwater Vehicles (UUVs), or even amphibious robots, which makes the acquisition of images from multiple domains costly.
@@ II-B Maritime and Marine Image Segmentation
Several efforts [16, 17, 18, 19] investigate sea-sky line inference and water obstacles detection for unmanned surface vehicles (USVs) using the Marine Semantic Segmentation Training Dataset (MaSTr1325), but do not explore the segmentation of virtual and real image regions, Vandael et al. [20], Muhadi et al. [21] and Wagner et al. [22] worked on automated river level monitoring. However, these methods do not utilize information across frames and may not generalize effectively for tasks near the water surface, especially when observing from beneath the water surface. We propose a method that can work above and beneath the water surface and utilizes temporal cues in consecutive frames. Unlike water segmentation at the water-air interface, real-virtual image segmentation prioritizes temporal motion and geometric details due to the robots’ first-person viewpoint, egomotion, and proximity to water surfaces.
@@ III-A AquaSim Simulator
$*{}^{*}$nly Blender and UE5 rendering engine supports ray-tracing with OptiX denoiser for photorealistic images;

### v2

@@ I INTRODUCTION
Advanced computer vision systems enable robots to comprehend scenes by analyzing pictorial information. Although this may work well in many cases, various real world scenarios can pose challenges due to the diverse nature of light propagation and scene complexity. The images received by human and machine visual sensors are often a complex superposition of direct imaging, reflections, refractions, scattering, and the likes. This complexity is particularly pronounced at the interfaces between different media. Consequently, the input to vision sensors contains real image regions formed by direct imaging and virtual areas typically formed by distortion, reflection, or refraction. Inherent complexities in differentiating between real and virtual regions pose challenges for both low-level and high-level computer vision tasks (see Fig. 1), such as depth and pose estimations, visual odometry, object recognition, semantic segmentation, etc. Therefore, real-virtual image segmentation to localize visually credible areas is crucial for marine robots, particularly in multimedia scenarios.
@@
Novel deep learning-based methods have significantly improved the performance of numerous computer vision tasks such as object recognition [1] and detection [2], depth estimation [3] and 3D navigation [4, 5, 6]. However, our investigation indicates that real-virtual image segmentation in multimedia scenarios remains a relatively underexplored area. One primary challenge is the scarcity of publicly available datasets for real-virtual image segmentation in multimedia scenarios. Additionally, existing underwater simulators often overlook the intricate modeling of water-air interface imaging, making synthetic data extremely difficult to acquire. At the same time, data-driven methods are heavily relying on large-scale datasets. Secondly, the absence of an efficient annotation pipeline for the real-virtual image segmentation task, particularly for images with visual scene ambiguity, results in manual annotation being more laborious than common computer vision tasks. Moreover, achieving this task usually involves the use of Unmanned Aerial Vehicles (UAVs), Autonomous Surface Vehicles (ASVs), Unmanned Underwater Vehicles (UUVs), or even amphibious robots, which makes the acquisition of images from multiple domains costly.
@@ II-B Maritime and Marine Image Segmentation
Several efforts [18, 19] investigate sea-sky line inference and water obstacles detection for unmanned surface vehicles (USVs) using the Marine Semantic Segmentation Training Dataset (MaSTr1325), but do not explore the segmentation of virtual and real image regions, Vandael et al. [20], Muhadi et al. [21] and Wagner et al. [22] worked on automated river level monitoring. However, these methods do not utilize information across frames and may not generalize effectively for tasks near the water surface, especially when observing from beneath the water surface. We propose a method that can work above and beneath the water surface and utilizes temporal cues in consecutive frames. Unlike water segmentation at the water-air interface, real-virtual image segmentation prioritizes temporal motion and geometric details due to the robots’ first-person viewpoint, egomotion, and proximity to water surfaces.
@@ III-A AquaSim Simulator
Only Blender and UE5 rendering engine supports ray-tracing with OptiX denoiser for photorealistic images;